OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (2, 6), what quadratic function is created?

OpenStudy (anonymous):

f(x) = -one eighth (x - 2)2 - 2 f(x) = one sixteenth (x - 2)2 + 2 f(x) = one eighth (x - 2)2 - 2 f(x) = -one sixteenth (x + 2)2 - 2

OpenStudy (anonymous):

@ganeshie8 @iPwnBunnies @ChristopherToni